OpenStudy (anonymous):

A revenue function is given by R(x)=15+114x-3x^2, where x is the number of units produced and sold. Find the value of x for which the revenue is maximum.

OpenStudy (anonymous):

thats if the derivetive is zero

OpenStudy (anonymous):

this one is simple, just find the derivative of of the revenue function and then get the critical points

OpenStudy (anonymous):

\[R(x)\prime = -6x + 114\] \[x = 114/6\] x=19

OpenStudy (anonymous):

thanks